OpenStudy (anonymous):

The surface area of a rectangular prism is 442 square feet. The length is 11 feet, and the height is 8 feet. What is the width of the prism? 5 ft 8 ft 2 ft 7 ft

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

I know 8 is wrong

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

7

OpenStudy (paki):

let me solve... i am alone here... and questions are 8

OpenStudy (anonymous):

A=2(wl+hl+hw) Just fill in the formula. But, 7 is the right answer.

OpenStudy (anonymous):

l: 11 h: 8 w: 7

OpenStudy (anonymous):

2(7*11+8*11+8*7)

OpenStudy (anonymous):

7*11=77 8*11=88 8*7=56

OpenStudy (anonymous):

77+88+56=221 221*2=442

OpenStudy (paki):

yeah... 7 is the rite answer... @girl8r7y

OpenStudy (anonymous):

Refer to the Mathematica attachment.